# The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals

### v2

@@ 3.1. Ethics
Each participant received $1.20\$$ upon completing the survey. To determine this figure, we ran a test of 20 participants, who took 5m56s (median) to complete it. Thus, the actual compensation was $12\$$/hr, which aligns with Prolific recommendations. For the signup survey of the cognitive interview, we set the study name to “Sign Up for Paid Interview on Data Donation” with a payment of $1.25\$$ ($13.68\$$/hr based on a loose completion estimate of 5 minutes). However, most respondents refused our interview request. Thus, in order to encourage participants interested in attending the interview to complete the screener, we lowered the compensation for the screening survey to $9.96\$$/hr. The interview reward is $15\$$, and all interviews lasted at most thirty minutes.
The survey is conducted through Qualtrics (Qualtrics, 2020), with the IP address collection feature disabled to protect respondents’ privacy. Interviews are conducted via Zoom (zoom, 2024), and participants are instructed to avoid using their real names when joining the meeting to ensure anonymity. Research participants are anonymously recruited through Prolific (Prolific, 2024), as described in greater detail below.
3.2. Survey Structure
@@
Figure 1. Survey design. Each participant is presented with a donation scenario and a survey introduction with two variations: recipient entity and disease. statements (PG(1)–PG(4) and AG(1)–AG(2)) are also presented in the donation scenario. Participants are then asked five sets of questions (survey questions). When the control statements are presented, the trust (dashed) is omitted. The willingness to donate and trust questions (dark-shaded) are associated with an open-text question. The scenario understanding is used as a filter. Participants who did not understand the donation scenario were excluded from the analysis. A sample screen that participants would encounter is provided in Figure 6.
@@
Our survey design is shown in Figure 1, and the full survey can be found in Appendix A. First, we present to the respondents a suvery introduction, which we detail in Figure 2.
@@
Figure 2. Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease. The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) (for Medicare and Services, 2024).
Then, we present to the respondents either a control (no statement) or an experimental statement. Experimental statements are composed of either one privacy statement, alone, or one privacy statement and one auditing statement. Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements. We have written these statements to align with PG(1)–PG(4) (see § 1), and we list them in Figure 3. We further distinguish the auditing guarantee into two different auditing statements to align with AG(1)–AG(2), and we list them in Figure 4.
PG(1): Anonymization — personally identifiable data will not be collected.
The privacy-preserving technology removes any personal identifiable information at the time of data collection, so that the data stored by the recipient entity is not linkable to its data owner (i.e., the data is anonymous).
PG(2): Access control — only relevant scientists can read data.
The privacy-preserving technology restricts data access to the authorized scientists (within the recipient entity) which are working on the treatment for the given disease.
PG(3): Data expiration — data will be deleted later.
The privacy-preserving technology discards the donated data, or makes it inaccessible after a given expiration time, which can be chosen by the data donor in the data collection agreement.
PG(4): Purpose restriction — data will not be used for other purposes.
The privacy-preserving technology ensures that the recipient entity can only use the donated data to develop a treatment for the given disease, and not in the context of any other research they may be working on (e.g., different disease).
Baseline: Control
No privacy-preserving guarantee is at all mentioned.
Figure 3. Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
AG(1): Expert auditing — the implementation of PET will be verified by experts.
An external advisory board of scientists and software engineers appointed by the recipient entity will regularly verify that the privacy-preserving technology is working as described. The results of this verification will be made public.
AG(2): Self-auditing — the implementation of PET can be verified by anyone.
Anyone interested, including the respondent and the experts the respondent trusts, will be able to verify that the privacy-preserving technology is working as described. Anyone, including the respondent and the experts the respondent trusts, can make their verification results public.
Baseline: Control
No auditing is at all mentioned.
Figure 4. Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
@@
  • scenario understanding, used to filter out respondents that report not understanding the scenario (see Section 3.4): “How would you rate your understanding of the above scenario?” (4 point Likert scale: Fully understand - Not understand).
@@
  • privacy expectations regarding the specific privacy guarantees we investigate, measured by their agreement with the statements presented in Figure 5 using the same Likert scale as willingness to donate. [^1] This question included an additional attention check with a sixth statement: “If I donate my data, I will meet Albert Einstein.” Respondents who did not answer ‘Very Unlikely’ were removed from the experiment, as detailed in Section 3.4. We assess participants’ expectations about all guarantees, regardless of what statement they were presented.
@@
  • demographics and experiences, including age, gender, education level, donation history, technical background (i.e., education background and job field relationships with computer technologies), and egocentricity(i.e., whether respondents or their close relatives have the disease, as defined in (Grgić-Hlača et al., 2020; Plane et al., 2017)).
@@
The donated medical records will be used for another purpose without my consent
@@
Figure 5. To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
@@
A sample screen that participants would encounter is provided in Figure 6.
Imagine that a non-profit organization (e.g., university, non-profit research institute) wants to develop a new treatment for Hypertension (High blood pressure). They need medical data from people with and without Hypertension (High blood pressure) to develop the treatment. They ask you to donate your medical record to help develop the treatment. Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
The non-profit organization uses a privacy-preserving technology to ensure that your data is deleted after a set period of time. You will be able to choose from a list of time period options how long your data will be stored. Anyone interested, including you and experts you trust, will be able to verify that the privacy-preserving technology is working as described. Anyone, including you and experts you trust, can post their verification results publically.
How would you rate your understanding of the above scenario? Fully understand Mostly understand Partially understand Not understand
Figure 6. Sample survey screen that composed of three segments. The first segment presents the survey introduction, which varies the disease and the entity among all participants. The second segment includes the randomized privacy statement and auditing statement, with color coding applied to emphasize these elements. The third segment contains the survey question that participants are required to answer.
3.3. Survey Refinement
@@
The final survey was refined through multiple iterations. Before launching the survey, we prescreened 49 users on the Prolific online platform (Prolific, 2024) by asking them to complete a short demographics survey and indicate their willingness to participate in a cognitive interview. Then, we conducted 17 cognitive interviews to ensure participants understood the survey correctly. Each interview involved presenting the survey to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements. We additionally presented all statements to the interviewees and asked for their understanding and perception of each of them. We continued refining the survey through cognitive interviews until no further constructive feedback was received. In the final interview, we observe a clear understanding of the statements and the donation scenario. For example, the participant described anonymization (PG(1)) as ”eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked” and access control (PG(2)) as ”some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in.” Our final survey can be found in Appendix A.
3.4. Survey Sampling
@@
Following the recommendation of (Tang et al., 2022), we recruited participants through Prolific, gathering responses from 560 individuals. To ensure focused results, we restricted respondents to adults residing in the U.S. and requested a gender-balanced distribution. Our recruitment included 272 men, 275 women, 9 nonbinary individuals, and 4 participants who chose not to disclose their gender.
@@
Of these 560 respondents, we sequentially excluded the following from the analysis: 35 that submitted incomplete responses; 20 that failed our attention check (§ 3.2); and 11 that indicated they did not understand the survey scenario (using scenario understand). Thus, our final dataset comprises 494 respondents. To enhance statistical power, we limit the levels of age and education to binary groups based on the data distribution among our participants. We report their demographics in Table 1 and the number of respondents assigned to each condition in Table 2.
Description Category $n$
@@
Table 1. Participant demographics. We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis, except for “age”, which is treated numerically.
@@
Table 2. Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 7).
3.5. Analysis
We analyzed the open-text questions about willingness to donate and trust (see § 3.2) using inductive-thematic open coding. Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.2%) responses. Then, they composed a final codebook and double-coded all responses. Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) (McDonald et al., 2019).
@@ 3.5. Analysis
3.6. Limitations
Although the four privacy-preserving guarantees examined in this study (§ 1) cover a broad range of PETs, they are not fully comprehensive and do not capture the full complexity of PETs or real-world threats, which may involve privacy issues beyond the scope of our investigation. The randomly presented disease may also influence the results. Furthermore, our approach may not fully reflect the intricacies of real-world donation scenarios, where multiple guarantees often coexist. While such complexities lie outside the primary focus of our research, qualitative investigations have been conducted in prior work (Aitken et al., 2016; Mori, 2016), and quantitative exploration of these aspects remains an interesting direction for future research.
@@ 3.6. Limitations
Additionally, while our study employs vignette scenarios that effectively mirror real-world behaviors (Redmiles et al., 2017), there may still be discrepancies between participants’ expressed privacy concerns and their actual behaviors, a phenomenon commonly referred to as the privacy paradox (Barth and de Jong, 2017).
@@ 4.1. RQ1: Understanding of PGs
Next, we answer the first research question (RQ1, see § 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?”. We analyze and report the survey responses and evaluate the rationales for the responses.
4.1.1. Analysis Setup
@@
We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 3) and auditing statements (listed in Figure 4). In Figure 7, we show the privacy expectations of respondents who were shown different privacy statements. To assess the statistical significance of the descriptive quantitative results presented above, we used binomial logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations. We summarize results, which we distinguish for non-profit and for-profit entities, in Figure 8.
@@
Figure 7. Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario. The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented. The 3rd column (“overall”) reports results across both entities. For example, the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected their data to be anonymized. The agreements are binarized here to present the overall tendency.
@@
Figure 8. Odd ratios and 95% confidence intervals of presenting different privacy statements on the respondents’ privacy expectations toward the two kinds of recipient entities, for-profit and non-profit.
4.1.2. Response Distribution and Statistical Analysis Results

### v3

@@ III-A Ethics
Each participant received 1.20\$1.20\\$ upon completing the survey. To determine this figure, we ran a test of 20 participants, who took 5m56s (median) to complete it. Thus, the actual compensation was 12\$12\\$/hr, which aligns with Prolific recommendations. Interview participants were first recruited through a screening survey that collected demographic information and assessed willingness to participate in interviews. Initially, we compensated participants 1.25\$\\$ for completing this 5-minute screener (approximately 13.68\$\\$/hr), titled “Sign Up for Paid Interview on Data Donation.” However, this above-average compensation led to only 8% interview attendance, as many users were motivated primarily by the screening fee.
To better align incentives and improve participation rates, we adjusted the screener compensation to 0.83\$\\$ (approximately 9.96\$\\$/hr), which better reflected platform norms for screening surveys. This adjustment increased our interview attendance rate to 21%. Participants who completed the 30-minute interviews were additionally compensated 15\$\\$.
III-B Donation Scenarios
PG(1) Anonymization: The privacy-preserving technology removes any personal identifiable information at the time of data collection, so that the data stored by the recipient entity is not linkable to its data owner (i.e., the data is anonymous).
PG(2) Access control: The privacy-preserving technology restricts data access to the authorized scientists (within the recipient entity) which are working on the treatment for the given disease.
PG(3) Data expiration: The privacy-preserving technology discards the donated data, or makes it inaccessible after a given expiration time, which can be chosen by the data donor in the data collection agreement.
PG(4) Purpose restriction: The privacy-preserving technology ensures that the recipient entity can only use the donated data to develop a treatment for the given disease, and not in the context of any other research they may be working on (e.g., different disease).
Baseline: N/A: (No privacy guarantee is at all mentioned.)
Figure 1: Privacy statements PG(1)–PG(4). One of the five statements (including the empty baseline statement) is randomly presented in the donation scenario immediately after the survey introduction.
AG(1): Expert auditing: An external advisory board of scientists and software engineers appointed by the recipient entity will regularly verify that the privacy-preserving technology is working as described. The results of this verification will be made public.
AG(2): Self-auditing: Anyone interested, including the respondent and the experts the respondent trusts, will be able to verify that the privacy-preserving technology is working as described. Anyone, including the respondent and the experts the respondent trusts, can make their verification results public.
Baseline: N/A: (No auditing statement is at all mentioned.)
Figure 2: Auditing statements AG(1)–AG(2). One of the three statements (including the empty baseline statement) is randomly presented in the donation scenario after a non-control privacy statement.
Many PETs involve complex mechanisms that cannot be adequately explained in brief consent forms without compromising informed consent principles. Additionally, consent forms must satisfy legal and ethical requirements while remaining accessible, as participants typically spend just minutes reviewing them [90, 91, 92].
To address this challenge, we align with real-world consent practices [42, 41] by measuring reactions to the high-level guarantees that PETs provide rather than explaining their technical implementations. This approach allows us to investigate fundamental privacy guarantees that appear consistently across different technologies—guarantees that are both meaningful to users and reflective of actual implementations:
  1. PG(1):
    Anonymization [93, 94]: data is not linkable to its owner (as defined in [95]);
  2. PG(2):
    Access control [4, 96, 97, 98]: data is accessible only by authorized people, specified in the data collection agreement;
  3. PG(3):
    Data expiration [99, 100, 101]: data is discarded or inaccessible after a given expiration time, specified in the data collection agreement;
  4. PG(4):
    Purpose restriction [24, 102]: data is only used for the stated collection purpose, specified in the data collection agreement;
In practice, implementing all guarantees simultaneously can be costly and challenging for real-world systems [43, 44]. This constraint is reflected in many consent frameworks, which often focus on providing a single guarantee—such as anonymization [42] or access control [41]. Following this practical approach, our study investigates the impact of each guarantee individually, rather than presenting them in combination. This targeted analysis enables us to identify which specific protections most significantly influence privacy expectations and, consequently, reported data-sharing intentions.
Beyond these privacy guarantees, the technical community has developed methods to verify that PETs function as intended. We also investigate how these auditing mechanisms affect user trust and willingness to donate data, specifically examining two approaches [^1] We considered but excluded government auditing as a separate category due to overlap with “expert auditing” and potential political bias (see § II).:
  1. AG(1):
    Expert auditing [103, 104]: engaging an aggregator-selected external advisory board to audit the system to verify that the PET functions as described in the data collection agreement;
  2. AG(2):
    Self auditing [105, 106]: granting anyone, including donors or external advisors appointed by them, the ability to perform such audits.
We present all statements in plain text format, avoiding metaphors and visual aids due to their mixed effectiveness in privacy communications [107, 108] and to align with standard practice in medical and survey consent forms [41, 42]. Figure 1 and Figure 2 show the exact text presented (e.g., the factor levels).
Justification on selecting the included PETs is in § -D.
III-C Survey Structure
@@
Figure 3: Survey design. Each participant is presented with a donation scenario and a survey introduction with two variations: recipient entity and disease. statements (PG(1)–PG(4) and AG(1)–AG(2)) are also presented in the donation scenario. Participants are then asked five sets of questions (survey questions). When the control statements are presented, the trust (dashed) is omitted. The willingness to donate and trust questions (dark-shaded) are associated with an open-text question. The scenario understanding is used as a filter. Participants who did not understand the donation scenario were excluded from the analysis. A sample screen that participants would encounter is provided in Figure 6.
@@
Our survey design is shown in Figure 3, and the full survey can be found in § -A. First, we present to the respondents a suvery introduction, which we detail in Figure 4.
@@
Figure 4: Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease. The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) [109].
Then, we present to the respondents either a control (no statement) or an experimental statement. Experimental statements are composed of either one privacy statement, alone, or one privacy statement and one auditing statement. Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements. Specifically, one statement from Figure 1 and one from Figure 2 are presented to each respondent.
@@
  • scenario understanding, used to filter out respondents that report not understanding the scenario (see § III-E): “How would you rate your understanding of the above scenario?” (4 point Likert scale: Fully understand - Not understand).
@@
  • privacy expectations regarding the specific privacy guarantees we investigate, measured by their agreement with the statements presented in Figure 5 using the same Likert scale as willingness to donate. [^2] This question included an additional attention check with a sixth statement: “If I donate my data, I will meet Albert Einstein.” Respondents who did not answer ‘Very Unlikely’ were removed from the experiment, as detailed in § III-E. We assess participants’ expectations about all guarantees, regardless of what statement they were presented.
@@
  • demographics and experiences, including age, gender, education level, donation history, technical background (i.e., education background and job field relationships with computer technologies), and egocentricity(i.e., whether respondents or their close relatives have the disease, as defined in [110, 111]).
@@
Figure 5: To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
Figure 6: Screenshot of the survey, which consists of three segments. The first segment presents the survey introduction, varying the disease and the entity across participants. The second segment presents the randomized privacy statement and auditing statement, highlighted with color coding for emphasis. The third segment poses the survey question that participants are required to answer.
@@
A screenshot of the survey is presented in Figure 6 to illustrate the survey design.
III-D Survey Refinement
We refine our survey thourgh a pilot study and multiple iterations of cognitive interviews, following the design of [90]. Inividuals are only allowed to participate in one of the studies.
We use the pilot study to test the survey design and the donation scenario. We found, for example, that no significant relationship between different diseases and donation willingness($p=0.32$, $\chi^{2}=17.00$) in the pilot survey ($N=81$ in the control condition). We thus present five common chronic diseases in the final survey (see § -A) to ensure participants can relate to the donation scenario and catch positive egocentricity measurement. We list the major takeaways in § -C.
We conducted 17 cognitive interviews with participants recruited through Prolific [89], selected from 49 users who completed our screening survey and indicated their willingness to participate in the interview. These interviews evaluated participants’ understanding of the survey content and helped facilitate consistent comprehension across respondents. Each interview involved presenting the survey to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements. We additionally presented all statements to the interviewees and asked for their understanding and perception of each of them.
@@ III-D Survey Refinement
We continued refining the survey through cognitive interviews until no further constructive feedback was received. In the final interview, we observed generally good understanding of the statements and the donation scenario. For example, the participant described anonymization (PG(1)) as “eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked” and access control (PG(2)) as “some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in.”
III-E Survey Sampling
@@
Following the recommendation of [112], we recruited participants through Prolific [89] and collected 560 responses. To ensure focused results, we restricted respondents to adults residing in the U.S. and requested a gender-balanced distribution. Our recruitment included 272 men, 275 women, 9 nonbinary individuals, and 4 participants who chose not to disclose.
@@
Of these 560 respondents, we sequentially excluded the following from the analysis: 35 that submitted incomplete responses; 20 that failed our attention check (§ III-C); and 11 that indicated they did not understand the survey scenario (using scenario understand). Thus, our final dataset comprises 494 respondents. To enhance statistical power, we limit the levels of age and education to binary groups based on the data distribution among our participants. We report their demographics in Table I and the number of respondents assigned to each condition in Table II.
TABLE I: Participant demographics. We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis, except for “age”, which is treated numerically.
Description Category $n$ %
@@
TABLE II: Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 7).
@@
III-F Analysis
We analyzed the open-text questions about willingness to donate and trust (see § III-C) using inductive-thematic open coding. Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.2%) responses. Then, they composed a final codebook and double-coded all responses. Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) [113]. Ultimately, six responses did not fit this scheme and classified as ‘Other’. The codebook of the open-text responses is presented in § -B.
@@ III-F Analysis
III-G Limitations
Although the four privacy-preserving guarantees examined in this study (§ I) cover a broad range of PETs, they are not fully comprehensive and do not capture the full complexity of PETs or real-world threats, which may involve privacy issues beyond the scope of our investigation. Furthermore, our approach may not fully reflect the intricacies of real-world donation scenarios, where multiple guarantees may coexist. While such complexities lie outside the primary focus of our research, qualitative investigations have been conducted in prior work [2, 54], and quantitative exploration of these aspects remains an interesting direction for future research.
@@ III-G Limitations
Additionally, our study relies on self-reported data, which is subject to well-documented biases, known as the privacy paradox, where individuals’ expressed privacy concerns often differ from their actual behaviors [114]. To address these limitations, we implemented several methodological safeguards. Firstly, we employed vignette-based scenarios, which research has shown to effectively mirror real-world behaviors and reduce hypothetical bias [115]. Secondly, our work is informed by prior research from 2019 [116] and replicated in 2022 [117] that demonstrated crowdsourced samples, specifically from Prolific [117], well approximate the security and privacy behavior, expectations, and knowledge of the general population for US adults between 18-50 who have at least some college education. While there may still be discrepancies between participants’ expressed privacy concerns and their actual behaviors, research confirms that this overambitiousness is systematic [118, 92]. It establishes survey data as a reliable upper-bound approximation of real-world behavior.
@@ IV-A RQ1: Understanding of PGs
We first address RQ1 (see § I): “How well do people understand and expect privacy guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?” We analyze survey responses and their underlying rationales.
IV-A1 Analysis Setup
@@
We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 1) and auditing statements (listed in Figure 2). In Figure 7, we show the privacy expectations of respondents who were shown different privacy statements. To assess the statistical significance of the descriptive quantitative results presented above, we used logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations. We summarize results, which we distinguish for non-profit and for-profit entities, in Figure 8.
@@
Figure 7: Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario. The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented. The 3rd column (“overall”) reports results across both entities. For example, the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected their data to be anonymized. The agreements are binarized here to present the overall tendency.
@@
Figure 8: Odd ratios and 95% confidence intervals of presenting different privacy statements on the respondents’ privacy expectations toward the two kinds of recipient entities, for-profit and non-profit.
IV-A2 Response Distribution and Statistical Analysis Results